OpenStudy (anonymous):

Hey what does the little circle mean in between a problem? I can't find that symbol anywhere! Please help!

OpenStudy (anonymous):

It means multiply

OpenStudy (anonymous):

(f * g)(x), but it's a circle, not a star

OpenStudy (anonymous):

Yea thats what I thought but I didnt excatly know what she was talking about

OpenStudy (anonymous):

an actual circle

OpenStudy (anonymous):

Oh okay never mind

OpenStudy (anonymous):

Hmmm well lets see.....

OpenStudy (anonymous):

is it a circle or a period in the middle

OpenStudy (anonymous):

Stupid pre-calc!

OpenStudy (anonymous):

Im not a big fan of math either.......

OpenStudy (anonymous):

circle just like the star, but a damn circle for some reason! lol

OpenStudy (anonymous):

Hmm....... I gotta think about this..... Lol

OpenStudy (anonymous):

Im assuming its multiplacation but Idk........ Because a multiplacation symbol has several ways to do it...... That looks like one of them

OpenStudy (anonymous):

Im not for sure...

OpenStudy (anonymous):

But, sometimes the dot=multiply

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

No. It means f composed of g.

OpenStudy (anonymous):

\((f\circ g)(x) = f(g(x))\)

OpenStudy (anonymous):

Ohh okay that makes sence now

OpenStudy (anonymous):

Crap! ok. thanks........somebody shoot me now.

OpenStudy (anonymous):

So for example: \[f(a) = a^2\] \[g(x) = x+3\] \[(f\circ g)(x) = f(g(x)) = (x+3)^2\]

OpenStudy (anonymous):

Thats exactly how I feel -_- 5 more math pages to do...

OpenStudy (anonymous):

Me too , -_-

OpenStudy (anonymous):

Its all review from way long ago I dont remember a dang thing!